40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 26 17:07:41 2019, gautam, Update, LSC, POX / POY calibration POX_POY_sensorNoise.pdfALSnoise_20191125.pdf
    Reply  Tue Nov 26 18:16:08 2019, rana, Update, LSC, POX / POY calibration 
       Reply  Wed Nov 27 12:16:52 2019, gautam, Update, LSC, ITMX and ITMY OSEMs with low and high circulating power ITMXshadowSensors.pdfITMYshadowSensors.pdf
          Reply  Wed Nov 27 13:14:02 2019, rana, Update, LSC, ITMX and ITMY OSEMs with low and high circulating power 
             Reply  Wed Nov 27 16:10:29 2019, gautam, Update, LSC, AOM reconnected 
Message ID: 15051     Entry time: Wed Nov 27 12:16:52 2019     In reply to: 15050     Reply to this: 15052
Author: gautam 
Type: Update 
Category: LSC 
Subject: ITMX and ITMY OSEMs with low and high circulating power 


The ITMX OSEMs report elevated noise in the 10-100 Hz band when we have high circulating power in the arm cavities, see Attachment #1. Since there is no LSC actuation on the ITMs in this state, this could be a radiation presssure effect, or could be scattered 1064nm light entering the OSEMs. The Oplevs don't report any elevated noise however. ITMY has the OSEM whitening broken for two channels, but the other two channels don't report as significant an increase as ITMX, see Attachment #2. I can't find the status of which OSEMs have the 1064nm blocking filters installed. The local damping loops are rolled off by ~100dB at 30 Hz, so the sensing noise re-injection should be attenuated by this factor, so maybe the OSEM sensor noise isn't the likely culprit. But radiation pressure didn't worsen the length noise in the past, even after our mirror cleaning and the increased PRG.


...maybe the opto-mechanical CARM plant is changing as a function of the CARM offset...

Attachment 1: ITMXshadowSensors.pdf  301 kB  | Hide | Hide all
Attachment 2: ITMYshadowSensors.pdf  289 kB  | Hide | Hide all
ELOG V3.1.3-